OpenStudy (anonymous):

Suppose a triangle has sides a, b, and c, and that a^2 + b^2 < c^2. Let be the measure of the angle opposite the side of length c. Which of the following must be true? ***choose ALL that apply!!*** A. cos theta < 0 B. cos theta > 0 C. The triangle is not a right triangle. D. a^2 + b^2 - c^2 = 2abcos theta ***my answer: C and A... What do u think @terenzreignz? :)

OpenStudy (anonymous):

since \[c^2\ge b^2+a^2\] \[c^2=a^2+b^2+\text{something}\]

terenzreignz (terenzreignz):

Well obviously a and b cannot both be true, right? Only one of them...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah haha :P well i think that this is an acute triangle.. am i right when i think that? :)

terenzreignz (terenzreignz):

Start with this...a\[\large a^2 + b^2 < c^2 \quad \rightarrow\quad a^2+b^2-c^2 < 0 \]

terenzreignz (terenzreignz):

So \[\large a^2+b^2-c^2 \] is negative okay?

OpenStudy (anonymous):

so this something is has to be negetive in order for the quanity to be plus so \[\cos\theta \le0\]

OpenStudy (anonymous):

ok :) how can i verify if it'll work? :D

terenzreignz (terenzreignz):

Let's be formal :D Since \(\large a^2+b^2-c^2\) is negative, and 2ab is positive, then \[\large \cos \theta \] which the law of cosines dictates to be equal to \[\large \frac{a^2+b^2-c^2}{2ab}\] should be... (positive/negative) ?

OpenStudy (anonymous):

negative?

terenzreignz (terenzreignz):

That's right :D So A is one of them...

OpenStudy (anonymous):

\[\cos \theta=\frac{ a^2+b^2-c^2 }{ 2ab } \le 0\]

terenzreignz (terenzreignz):

positive @iheartfood ? It has a negative numerator and a positive denominator -.-

OpenStudy (anonymous):

okay yay! :) so it would be A and C then? since this isn't a right triangle? :D if so, what about D?! :D

terenzreignz (terenzreignz):

Can't be a right triangle since cos theta is not 0... now... as for D...

terenzreignz (terenzreignz):

Doesn't that like follow directly from the law of cosines? \[\Large c^2 = a^2 + b^2 -2ab \ \cos(\theta)\]

OpenStudy (anonymous):

yeah haha sorry... totally over thought it and over thought it wrongly ahha.. put yeah i see that its negative @terenzreignz cuz of that!

terenzreignz (terenzreignz):

You can just bring \(c^2\) to the other side and do the same to 2ab cos(theta)

terenzreignz (terenzreignz):

\[\Large c^2\color{red}{-c^2}\color{green}{+2ab \ \cos (\theta)} = a^2 + b^2 -2ab \ \cos(\theta)\color{red}{-c^2}\color{green}{+2ab \ \cos(\theta)}\]

OpenStudy (anonymous):

okay! so it's A and C then? :) B and D just aren't the answers here?!

terenzreignz (terenzreignz):

D is just the law of Cosines rearranged...

OpenStudy (anonymous):

so D is also an answer then? :O

terenzreignz (terenzreignz):

yes

OpenStudy (anonymous):

okay! so A,C, and D it is! :) thx!! :D

OpenStudy (anonymous):

okay! so A,C, and D it is! :) thx!! :D